Express 2x^2+2y^2-12x+16y=-32 in standard form for a circle. Identify the radius and center of the circle.
Answer by venugopalramana(3286)

Express 2x^2+2y^2-12x+16y=-32.DIVIDING BY 2
x^2+y^2-6x+8y=-16
[X^2-2(X)(3)+3^2]-3^2+[Y^2+2(Y)(4)+4^2]-4^2=-16
(X-3)^2+(Y+2)^2=9=3^2
THIS IS IN THE STD.FORM OF
(X-H)^2+(Y-K)^2=R^2
HENCE CENTRE = (H,K) = (3,-2)
RADIUS = R =3
in standard form for a circle. Identify the radius and center of the circle.